# Junhao Chen* 1122    Xiang Li* 33    Xiaojun Ye 1144    Chao Li 55 Zhaoxin Fan 66

### v1

@@ 4.3 Results
Fig. 4 showcase the evolution and refinement of 3D models at different Idea-2-3D stages and how the caption-based baseline works. Fig. 5 shows the iterative self-improvement process of a case in Idea-2-3D. These visual representations illustrate the framework’s effectiveness in aligning with user IDEAs and the incremental improvements achieved through its iterative process.
4.4 Ablation Study
In this section, we provide ablative user studies to shed light on the effectiveness of critical design choices in Idea-2-3D. Qualitative and quantitative results on shown in Fig. 6 and Table. 2 respectively. Note that in previous experiments, we use a default iteration cap of 5, but here this limit is removed so that the end of iterative refinement is decided solely by whether the LMM agent 5 is satisfied or not. As can be seen in the experiments (sixth iteration and seventh iteration) for the Polemon case in Fig. 6, removing critical modules (including memory, feedback and storing previous models in memory) lead to slow quality convergence while our full Idea-2-3D converges within 3 iterations.
Figure 6: Key module ablation across iterations. Note that the full model reaches satisfactory results (judged by the LMM agent 5) at iteration 3.
The quantitative ablative user study in Tab. 2 shows that: (1) Our proposed memory module can effectively prevent the quality of 3D model generation from divergence (deviating from the user input IDEA). (2) Our proposed LMM feedback agent can effectively improve the speed of quality convergence, i.e., the generated 3D model can well align with user input IDEAs in shorter iteration rounds. (3) If the storage information of the previous model in the memory module is removed, there will be a lower quality upper limit (whether models satisfy IDEAs) after convergence.
Table 2: Ablation study for Idea-2-3D components, collectively obtained at standard iteration rounds ($N=5$) using the 21 test samples presented in this paper.
Satisfying(%) iter 1 iter 2 iter 3 iter 4 iter 5
w/o memory 64.87 60.32 55.34 48.98 41.36
w/o feedback 77.83 80.41 82.67 84.03 85.32
w/o previous Model 78.31 84.58 85.87 85.41 84.49
full model 78.20 89.45 95.23 96.18 96.95
Quality comparisons with commercial models. Qualitative and quantitative comparisons with commercial models are presented in Fig. 7 and Table. 3 respectively, showing that our Idea-2-3D enables a new 3D AIGC capability that existing commercial models fail to demostrate.
Efficiency. Since generation from $N$ prompts can run in parallel, the inference speed is agnostic of $N$. Idea-2-3D generates a result in about 20-30 minutes, since the average iteration round is 2-3 and each iteration takes about 10 minutes. For reference, existing commercial methods typically generate a model in about 5-10 minutes, e.g., 3DTopia and One-2-3-45++ takes about 5 minutes on average, while Luma and Meshy take about 10 minutes. Our method is comparable to them in scale (tens of minutes).
Figure 7: Comparison with commercial T-2-3D models. Prompts are the same as Fig. 2.
Table 3: Quantitative comparison with commercial T-2-3D models.
Method CLIP Score normal-$\uparrow$ Which is best normal-$\uparrow$ Satisfying normal-$\uparrow$
3DTopia 0.2512 2.2% 4.5%
One-2-3-45++ 0.2805 4.3% 24.7%
Luma 0.3068 8.7% 43.2%
Meshy 0.2827 6.5% 32.3%
Ours 0.3611 78.3% 94.2%
5 Conclusion
Idea-2-3D, utilizing an LMM agent collaboration framework, revolutionizes 3D AIGC by automating the creation of models from high-level, interleaved multimodal user inputs (IDEAs). This innovative system excels in integrating text, images, and 3D models, underpinned by a unique iterative process that enhances model coherence and visual alignment with IDEAs. User studies underscore its superiority in satisfaction and comparative quality, marking Idea-2-3D as a significant advancement in 3D AIGC and a benchmark for future design tools.

### v2

@@ Appendix D Appendix: Visualization of Self-iterative Refinement
Fig. 9 showcase the evolution and refinement of 3D models at different Idea23D stages and how the caption-based baseline works. Fig. 10 shows the iterative self-improvement process of a case in Idea23D.
Figure 9: Comparison between prompts for "Caption-based T-2-I prompt for initial round", "Idea2-3-D prompt for initial round" and "Iterative self-refinement Idea23D end round" and the comparison between generated 3D models.
Figure 10: An example of the process for a 3D model generated through the Idea23D framework. An initial prompt is first generated based on the input IDEA and $p_{\text{gen}}$, then multiple images are generated from the T-2-I model, and then the corresponding draft 3D model is generated for each image separately. The most appropriate 3D model is selected based on $p_{\text{select}}$ and feedback is generated. The results are output after several rounds of iterations. Green rectangles indicate text prompts for T-2-3D. blue rectangles indicate: modification suggestions given by agent based on draft 3D models and $p_{\text{fb}}$.
Appendix E Appendix: Efficiency
Since generation from $N$ prompts can run in parallel, the inference speed is agnostic of $N$. Idea23D inference speed mainly depends on the speed of the T-2-I and I-2-3D models, as the average number of iterations is 2-3. For example, with the Idea23D implemented using GPT-4V + zero123 + DALLE, generating an optimal result takes about 10 minutes (zero123 requires approximately 4 minutes to generate a 3D model from an image). In the Colab implementation mentioned in our abstract, we use GPT-4V + TripoSR + SDXL to implement the Idea23D framework, and it takes about 5 minutes to generate a final 3D model (with 3 iterations). For reference, existing commercial methods typically generate a model in about 5-10 minutes. For example, 3DTopia  [^3] https://github.com/3DTopia/3DTopia and One-2-3-45++Liu et al. (2023c) take about 5 minutes on average, while Luma [^4] https://lumalabs.ai/genie?view=create and Meshy  [^5] https://www.meshy.ai/ take about 10 minutes. Recent advanced methods such as TripoSR Tochilkin et al. (2024), InstantMesh Xu et al. (2024b), and LGM Tang et al. (2024) have compressed the time to generate 3D models from images to within 1 minutes.
Appendix F Appendix: More Specific Domains
We tested our approach within the domain of automated design and modeling of car, chair, and cloth. The four views are plotted as a result of 3D modeling, and the text prompts above indicate the User Idea we entered. The result is shown in Fig. 11. The abstract IDEA input is above the picture.
Figure 11: Domain-specific results of Idea23D.